I need some help on this problems please!
Solve the system with the addition method: 5 x 4y 3 20x 16y 12 Answer: ( )
Solve the following system of equations with the substitution method: 4y +3 l-4x +16y -12 Answer: ( ) Enter your answers as integers or as reduced fractions in the form A/B. If there is no solution, t there are an infinite number of solutions, specify their form in the blanks in terms ofy
Solve the system with the addition method: 5 x 4y 3 20x 16y 12 Answer: ( )
You have one type of chocolate that sells for $2.00/lb and another type of chocolate that sells for $4.20/lb. You would like to have 17.6 lbs of a chocolate mixture that sells for $3.00/lb. How much of each chocolate will you need to obtain the desired mixture? You will need Ibs of the cheaper chocolate and Ibs of the expensive chocolate.
Bryce invests $3100 in two different accounts. The first account paid 8 %, the second account paid 7 % in interest. At the end of the first year he had earned $236 in interest. How much was in each account? at 8 % at 7 %
Answer #1

4y+3 -20% -164=t2 -4/44+ 3) +16y =-12 J164- 12+ 164 =-12 ho lds for all it 20%+16y = -12 of ,y valve Crtinite Solulons -20% -